OpenStudy (anonymous):

Given the function defined by g(x)=5x-3, find g(1/5) Ok so I worked this out and got 7/5 but my mathlab marked it wrong, im pretty sure it is 7/5 no?

OpenStudy (blurbendy):

If g(x) is defined as being equal to 5x - 3, then put 1/5 in for x. 5 ( 1 / 5) -3 = 5 /5 - 3 = 1 - 3 = -2

OpenStudy (anonymous):

-2 ?